Not sure about the category as I am no math wiz in any way shape or form, but here's the equation going around and I'm in a debate about it.

0+50x1-60-60x0+10=???

The answer I get is 10 and others are saying it's 0. I came up with the number in my head and then using a calculator I got the same answer by entering the equation exactly as it's written. My 14 year old son read the equation and instantly said 10. Which is the correct answer?

This is how I came to my answer of 10
0+50=50
50x1=50
50-60=-10
-10-60=-70
-70x0=0
0+10=10
1
Expert's answer
2013-02-19T10:09:01-0500
From school math we know that first you must multiply and than sum .

1) 50x1=50
2) 0+50x1 = 0 + 50 = 50
3) 50 - 60 = -10
4) 60x0= 0
5)-10 - 0 = -10
6) -10 + 10 = 0

the answer is 0